OpenStudy (anonymous):

Given: sin(x) = 2/3 and cos(y) = 3/4 Find the values of: a) sin(x+y) b) cos(x-y) c) tan(x+y)

OpenStudy (anonymous):

They are the general formula you have to learn: \[\sin(x+y) = \sin(x) \cos(y) + \cos(x) \sin(y)\] \[\cos(x-y) = \cos(x)\cos(y) + \sin(x)\sin(y)\] \[\tan(x+y) = \frac{\tan(x) + \tan(y)}{1 - \tan(x)\tan(y)}\]